### v1

# Quantum Metrology via Floquet-Engineered Two-axis Twisting and Turn Dynamics

Abstract
The core of quantum metrology lies in utilizing entanglement to enhance measurement precision beyond standard quantum limit. Here, we utilize the Floquet-engineered two-axis twisting (TAT) and turn dynamics to generate non-Gaussian states for quantum metrology. By employing both analytically semi-classical and quantum approaches, we find that the desired $N$-particle non-Gaussian state can be produced within a remarkably short time $t_{\mathrm{opt}}\propto\ln{N}/{N}$, and its quantum Fisher information $F^{\mathrm{opt}}_{\mathrm{Q}}\propto N^{2}$ approaches the Heisenberg limit. Moreover, using the Floquet-engineered anti-TAT-and-turn, we may implement an efficient interaction-based readout protocol to extract the signal encoded in this non-Gaussian state. This Floquet-engineered anti-TAT-and-turn approach offers a viable method to achieve effective time-reversal dynamics for improving measurement precision and resilience against detection noise, all without the need to invert the sign of the nonlinear interaction. This study paves the way for achieving entanglement-enhanced quantum metrology via rapid generation of cat-like states at high particle numbers through continuous Floquet engineering.
Introduction.— Quantum metrology utilizes quantum entanglement to enhance measurement precision from the standard quantum limit (SQL) to the Heisenberg limit (HL) [1, 2, 3, 4, 5, 6, 7, 8], offering potential applications in atomic clocks [9, 10, 11, 12, 13, 14, 15, 16], magnetometers [17, 18, 19, 20, 21, 22, 23, 24], gyroscopes [25, 26, 27, 28], gravimeters [29, 30, 31] and other sensors. The ultimate measurement precision of an estimated phase $\phi$ is determined by the quantum Cramér-Rao bound (QCRB) $\Delta\phi\geq 1/\sqrt{F_{Q}}$, defined by the quantum Fisher information (QFI) ${F_{Q}}$  [32, 33, 34]. How to prepare an achievable entangled state with high QFI is a top priority. One-axis twisting (OAT) dynamics [35, 36] stands out as a crucial mechanism for generating entanglement. Using OAT dynamics, the Gaussian-shaped spin squeezed states have been successfully demonstrated in several proof-of-principle experiments [37, 38, 39, 40, 41, 42]. Although non-Gaussian entangled states [43, 44, 45, 46] (especially the spin cat states [47, 48, 49] or Greenberger-Horne-Zeilinger (GHZ) state [50, 51]) offer higher QFI [52, 53], they require significantly longer evolution times for preparation, presenting a substantial practical challenge.
It has been demonstrated that OAT-and-turn [54, 55] and two-axis twisting (TAT) [56, 57, 58] can achieve faster entanglement generation. Although TAT exhibits a superior speedup, its interaction form does not naturally exist in known physical systems [59, 60]. In principle, an effective TAT interaction can be realized with OAT-and-turn, in which one can transform OAT into TAT by applying suitable transverse coupling [61, 62, 63, 64]. The coupling field can be a train of periodic pulses [61, 62], a periodically modulated continuous field [63], or even a sequence designed by machine learning [64]. Since the turn dynamics can accelerate OAT to achieve fast entanglement generation, a natural question arises: can it accelerate TAT to achieve faster entanglement generation?
On the other hand, extracting the signal encoded in entangled states generally necessitates single-particle-resolution measurements of parity [51, 65], high-order observables [66, 67], or full probability distribution [44, 45]. However, the single-particle-resolution measurements are very susceptible to detection noise [68, 69]. To address this challenge, time-reversal interaction-based readout provides a powerful protocol for achieving high-precision Heisenberg-limited measurements [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]. This requires to reverse the dynamics of an interacting many-body quantum system, which is usually accomplished by inverting the sign of nonlinear interaction [73, 74, 75]. Can one realize time-reversal TAT-and-turn dynamics to achieve Heisenberg-limited measurement without changing the sign of nonlinear interaction?
In this Letter, we employ Floquet-engineered TAT-and-turn dynamics to efficiently create non-Gaussian entangled states and carry out the necessary time-reversal readout. By applying a suitable transverse coupling field, an effective TAT-and-turn dynamics of an ensemble of Bose condensed two-level atoms can be realized via Floquet engineering. Using both analytically semi-classical and quantum treatments, we find that non-Gaussian entangled states with high QFI can be generated in much shorter time. Furthermore, the time-reversal readout can be realized by applying another Floquet-engineered TAT-and-turn dynamics without flipping the sign of nonlinear interaction. It allows to efficiently extract the signal encoded in the prepared non-Gaussian entangled states and thus provides a robust approach for enhancing measurement precision. Our work develops an efficient protocol for performing entanglement-enhanced quantum metrology via rapid generation and detection of cat-like states through continuous Floquet engineering.
[…]
$\hat{H}_{\mathrm{FE}}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos(\omega t)\hat{J}_{\alpha},$ (1)
[…]
Figure 1: (a) An ensemble of two-level particles ($\chi$ denoting the nonlinear interaction between particles) coupled via an external coupling field with detuning $\delta$ and (b) periodically modulated Rabi frequency $\Omega(t)=\Omega_{0}\cos(\omega t)$. (c) Schematic diagram of Floquet-engineered TAT-and-turn. The classical phase-space trajectory for (i) one-axis-twist $\chi\hat{J}_{z}^{2}$, (ii) rotation induced by energy imbalance $\delta\hat{J}_{z}$, (iii) rotation induced by modulated linear coupling $\Omega_{0}\cos(\omega t)\hat{J}_{\alpha}$. The combination of these three terms results in (iv) the effective TAT-and-turn dynamics.
Figure 2: The maximum QFI of states generated by the Floquet-engineered TAT-and-turn dynamics. (a) The maximum QFI $F_{Q}^{\mathrm{max}}$ versus evolution time $\chi t$ under different $\delta$ with particle number $N=100$. The amplitude and initial phase of Rabi frequency are chosen as $\Omega_{0}=2\pi\times 100N\chi$ and $\alpha=0$, respectively. (b) The time evolution of maximum QFI for Floquet-engineered TAT-and-turn with optimal detuning $\delta_{\mathrm{opt}}$, and the results of OAT, OAT-and-turn, TAT, and TAT-and-turn are presented for comparison. (c) The probability distribution and Husimi distribution (inset) of the optimal state with the largest QFI in (b), where the probability distribution is projected onto the eigenstates of $\hat{J}_{y}$ with $P_{m_{y}}=|\langle m_{y}|\psi\rangle|^{2}$. (d) The critical detuning $\delta_{\mathrm{opt}}$ and (e) the critical evolution time $\chi t_{\mathrm{opt}}$ corresponding to (f) the optimal QFI $F_{Q}^{\mathrm{opt}}$ for different particle numbers $N$, where the semi-classical predictions are $\delta_{\mathrm{opt}}^{\mathrm{SC}}/(N\chi)\approx 0.3135$, $t_{\mathrm{opt}}^{\mathrm{SC}}={3(1.9+0.55\ln{N})}/{N}$ and $(F_{Q}^{\mathrm{opt}})^{\mathrm{SC}}/N\approx 0.97N$, respectively. The fitted line for the optimal $F_{Q}^{\mathrm{max}}$ is $F_{Q}^{\mathrm{opt}}/N\simeq 0.77N$.
[…]
$\hat{H}_{I}$ $=$ $\hat{U}^{\dagger}\hat{H}_{\rm FE}\hat{U}-i\hat{U}^{\dagger}\partial_{t}\hat{U}$ (2)
$=$ $\chi(\cos\gamma\hat{J}_{z}+\sin\gamma\hat{J}_{\beta})^{2}+\delta(\cos\gamma\hat{J}_{z}+\sin\gamma\hat{J}_{\beta}).$
Then, we perform Fourier decomposition on this Hamiltonian. Using the Euler formula $e^{i\gamma}=\cos{\gamma}+i\sin{\gamma}$ and the Jacobi-Anger expansion $e^{iz\sin\theta}=\mathcal{{J}}_{0}(|z|)+\sum_{n=1}^{\infty}[\mathcal{{J}}_{n}(z)e^{in\theta}+(-1)^{n}\mathcal{{J}}_{n}(z)e^{-in\theta}]$ with $\mathcal{{J}}_{n}(z)$ being the $n$-th Bessel function of the first kind, the Hamiltonian (2) becomes
$\hat{H}_{I}=\hat{H}_{0}^{I}+\sum_{n=1}^{\infty}(\hat{H}_{n}^{I}e^{in\omega t}+\hat{H}_{-n}^{I}e^{-in\omega t}),$ (3)
where $\hat{H}_{0}^{I}=\frac{\chi}{2}[(1+L_{0})\hat{J}_{z}^{2}+(1-L_{0})\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z}$ and $\hat{H}_{n}^{I}$ represents $n$-order Hamiltonian with $L_{n}=\mathcal{{J}}_{n}\left(2\Omega_{0}/\omega\right)$ and $K_{n}=\mathcal{{J}}_{n}\left(\Omega_{0}/\omega\right)$ [88]. Finally, based on the Floquet-Magnus expansion [89, 90], up to the order of $\omega^{-1}$, the effective Hamiltonian for Eq. (11) can be written as $\hat{H}_{\mathrm{F}}^{\rm eff}=\hat{H}_{0}^{I}+\sum_{n=1}^{\infty}\frac{1}{n\omega}([\hat{H}_{n}^{I},\hat{H}_{-n}^{I}]-[\hat{H}_{n}^{I},\hat{H}_{0}^{I}]+[\hat{H}_{-n}^{I},\hat{H}_{0}^{I}])+\mathcal{O}(\omega^{-1})$. When $\omega$ is sufficiently large ($\omega\gg N\chi$), it reduces to $\hat{H}_{\mathrm{F}}^{\rm eff}\simeq\hat{H}_{0}^{I}$ [88]. For a conserved collective spin, $\hat{J}^{2}=\hat{J}_{\alpha}^{2}+\hat{J}_{\beta}^{2}+\hat{J}_{z}^{2}=\frac{N}{2}(\frac{N}{2}+1)$ is a constant, the effective time-independent Floquet Hamiltonian is equivalent to
$\hat{H}_{\rm F}=-\frac{\chi}{2}[(1+L_{0})\hat{J}_{\alpha}^{2}+2L_{0}\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z}.$ (4)
Setting the ratio $\Omega_{0}/\omega\simeq 1.6262$, we have $L_{0}=\mathcal{{J}}_{0}\left(2\Omega_{0}/\omega\right)=-1/3$, thus the Hamiltonian (4) becomes an effective TAT-and-turn Hamiltonian
$\hat{H}_{\rm TATNT}=\chi_{\rm eff}(\hat{J}_{\beta}^{2}-\hat{J}_{\alpha}^{2})+\delta_{\rm eff}\hat{J}_{z},$ (5)
with the effective interaction strength $\chi_{\mathrm{eff}}=\chi/3$ and the effective detuning $\delta_{\mathrm{eff}}=K_{0}\delta\simeq 0.404\delta$. Since the TAT-and-turn Hamiltonian (5) is derived from the Hamiltonian (9) with the Floquet theorem, we refer the original time-evolution to the Floquet-engineered TAT-and-turn (labelled as FE-TATNT) dynamics. To evaluate the Floquet-engineered TAT-and-turn dynamics, we also calculate the dynamics of the ideal TAT-and-turn Hamiltonian (5) for comparison.
Varying system parameters, the Floquet-engineered TAT-and-turn dynamics can be reduced to typical TAT, OAT-and-turn, or OAT dynamics. When $\delta=0$, the effective TAT-and-turn Hamiltonian (5) can be reduced to an effective TAT Hamiltonian [56, 57, 58] $\hat{H}_{\rm TAT}=\chi_{\rm eff}(\hat{J}_{\beta}^{2}-\hat{J}_{\alpha}^{2})$. Further, setting $\Omega(t)=\Omega_{0}$ as a constant and $\delta=0$, the Hamiltonian (9) is reduced to an OAT-and-turn Hamiltonian [54, 55] $\hat{H}_{\mathrm{OATNT}}=\chi\hat{J}_{z}^{2}+\Omega_{0}\hat{J}_{\alpha}$. If $\delta=0$ and $\Omega(t)=0$, the Hamiltonian (9) becomes an OAT Hamiltonian $\hat{H}_{\mathrm{OAT}}=\chi\hat{J}_{z}^{2}$.
Characterization of full metrological potential for a given state.— The metrologically useful entangled states created by the above different schemes typically correspond to different optimal sensing directions. To evaluate the full metrological usefulness of the states generated by different schemes, we first quantify the best achievable measurement precision using the maximum QFI $F_{Q}^{\rm max}$ for all potential sensing orientations. For an arbitrary state, one can determine a $3\times 3$ quantum Fisher information matrix (QFIM) $\mathcal{F}$  [91] with matrix elements $\mathcal{F}_{\mu\nu}=2\langle\{\hat{J}_{\mu},\hat{J}_{\nu}\}\rangle_{\psi}-4\langle\hat{J}_{\mu}\rangle_{\psi}\langle\hat{J}_{\nu}\rangle_{\psi}$, where $\mu,\nu=x,y,z$. By solving the eigenequation $\mathcal{F}\vec{n}=\lambda\vec{n}$, one can obtain three eigenvalues $\lambda_{i}$ and their corresponding eigenvectors $\vec{n}_{i}$ (where $i=\{1,2,3\}$). The maximum eigenvalue $\lambda_{\rm max}$ and its eigenvector $\vec{n}_{\rm max}=(n_{x}^{\rm max},n_{y}^{\rm max},n_{z}^{\rm max})^{\mathrm{T}}$ correspond to the maximum QFI $F_{Q}^{\rm max}$ and the optimal sensing direction (whose optimal generator is $\hat{J}_{\vec{n}_{\rm max}}=n_{x}^{\rm max}\hat{J}_{x}+n_{y}^{\rm max}\hat{J}_{y}+n_{z}^{\rm max}\hat{J}_{z}$), respectively. Therefore, one can use $F_{Q}^{\rm max}$ to determine the full metrological potential for quantum sensing and rotate the state until the optimal generator $\hat{J}_{\vec{n}_{\rm max}}$ matches the Hamiltonian for a given sensing purpose [91].
Fast generation of metrologically useful non-Gaussian entangled states with Floquet-engineered TAT-and-turn.— Using the above maximum QFI $F_{Q}^{\rm max}$, below we analyze the full metrological potential of the states generated by Floquet-engineered TAT-and-turn dynamics (9). Without loss of generality, we choose $\alpha=0$ hereafter. It is advantageous to begin with an initial spin coherent states on one of the poles along $z$-axis ($\lvert\uparrow\rangle^{\otimes N}$ or $\lvert\downarrow\rangle^{\otimes N}$), which correspond to two saddle points $(0,0,\pm 1)$ in the classical phase-space of TAT-and-turn when $\left\lvert\delta_{\rm eff}/\chi_{\rm eff}\right\lvert<N$, see Fig. 1 (c). To identify the optimal condition for Floquet-engineered TAT-and-turn, we analyze the dependence of $F_{Q}^{\rm max}$ on the detuning $\delta$ and the evolution time $\chi t$ in Fig. 2 (a). It clearly shows that the maximum QFI has a symmetrical distribution with respect to $\delta=0$. Given the particle number $N=100$, the maximum QFI reaches its optimal value $F_{Q}^{\mathrm{opt}}\simeq 0.7783N^{2}$ when the evolution time $\chi t_{\mathrm{opt}}\approx 0.132$ and the detuning $\delta_{\mathrm{opt}}\approx\pm 0.3135N\chi$. In Fig. 2 (b), we show the time evolution of $F_{Q}^{\rm max}$ for OAT, OAT-and-turn, TAT, TAT-and-turn ($\hat{H}_{\rm TATNT}$ with $\delta_{\mathrm{eff}}=K_{0}\delta_{\mathrm{opt}}$), and Floquet-engineered TAT-and-turn ($\hat{H}_{\rm FE}$ with $\delta=\delta_{\mathrm{opt}}$). The results of Floquet-engineered TAT-and-turn are well consistent with the ideal TAT-and-turn, thereby confirming the validity of the effective TAT-and-turn Hamiltonian (5). Compared with other schemes, the states generated via Floquet-engineered TAT-and-turn dynamics may achieve the largest $F_{Q}^{\rm max}$ when $0.10\lesssim\chi t\lesssim 0.18$. In terms of the best achievable QFI, the Floquet-engineered TAT-and-turn outperforms the OAT by approximately $54\%$, OAT-and-turn and TAT by around $20\%$.
Under the Floquet-engineered TAT-and-turn with $\chi t_{\mathrm{opt}}$ and $\delta_{\mathrm{opt}}$, the generated state is a non-Gaussian entangled state of two main parts distributed near the poles of the $y$-axis. Such a non-Gaussian state is similar to a spin cat state [47, 48, 49], see Fig. 2 (c). Moreover, the best achievable QFI of Floquet-engineered TAT-and-turn exhibits the Heisenberg scaling $F_{Q}^{\rm max}\propto N^{2}$, which is consistent with the QCRB of a spin cat state. Correspondingly, the evolution time for generating the state obeys $\chi t_{\rm opt}\propto\ln{N}/{N}$. This indicates that the required time can be very short for large particle number $N$, far less than that of the OAT dynamics, which is of essential importance and feasibility for overcoming decoherence.
The Heisenberg scaling, the critical detuning and the critical evolution time can be analytically found by using semi-classical treatment [54, 92]. Since there is a special symmetry in the classical phase space of the TAT-and-turn Hamiltonian (5), the states generated by Floquet-engineered TAT-and-turn dynamics always have symmetric probability distributions. Just like the resembled spin cat state in the inset of Fig. 2 (c), whose variance is approximately stretched along a certain axis, one can estimate the global optimal parameter $\delta_{\rm opt}^{\rm SC}=\pm(\sqrt{2}-1)N\chi/(3K_{0})\approx\pm 0.3135N\chi$ by the upper boundary of the QFI from the Heisenberg limit [34]. Due to the center of the uncertainty patch is fixed, the TAT-and-turn dynamics merely alters the distribution by stretching and squeezing. Thus, the relevant timescale can be estimated by the time interval between the boundary points of the uncertainty patch along the separatrix [92], which can be given as [88]
$\chi t_{\mathrm{opt}}^{\mathrm{SC}}\simeq\frac{3(1.9+0.55\ln{N})}{N}.$ (6)
Based on the above analysis, the best achievable QFI exhibits the Heisenberg scaling, which matches the ultimate precision bound of spin cat states [88]. As shown in Fig. 2 (d-f), the above results are in accordance with the numerical results of Floquet-engineered TAT-and-turn. According to Eq. (6), the Floquet-engineered TAT-and-turn dynamics exhibits significantly shorter timescale for large particle numbers compared to the OAT and it is even faster than the GHZ-like state generation with effective three-body interaction [93]. Thus our protocol is advantageous in mitigating decoherence in practical quantum metrology experiments, showing remarkable metrological usefulness.
Figure 3: (a) Measurement precision yielded from Floquet-engineered anti-TAT-and-turn. The results of time-reversal readout with OAT, OAT-and-turn, TAT and TAT-and-turn are presented for comparison. The QCRB of Floquet-engineered TAT-and-turn is also shown (green circle). (b) Robustness against detection noise. The metrological gain is defined as $G=20\log_{10}\left[{(\Delta\phi)_{\mathrm{SQL}}}/{\Delta\phi}\right]$ with $(\Delta\phi)_{\mathrm{SQL}}=1/{\sqrt{N}}$ denoting the measurement precision of SQL. Green dashed line is the result of optimal measurement $\hat{J}_{\vec{m}_{\rm max}}$ after Floquet-engineered anti-TAT-and-turn with $\chi t=0.12$. Green dashed-dot and yellow dot lines are the results with direct parity measurement for non-Gaussian state generated by Floquet-engineered TAT-and-turn at $\chi t=0.12$ and the GHZ state generated by OAT at $\chi t=\pi/2$, respectively. Here, the amplitude and initial phase of Rabi frequency are chosen as $\Omega_{0}=2\pi\times 50N^{2}\chi$ and $\alpha=0$ with particle number $N=100$ and the estimated phase $\phi=1/1000$.
Optimal measurement with Floquet-engineered anti-TAT-and-turn.— It is crucial to approach the ultimate precision bound (i.e. the QCRB) set by the QFI. For a tiny estimated phase, time-reversal readout may approach the QCRB without single-particle resolved detection [69, 70, 71, 72, 68, 73, 74, 75, 76, 77, 78, 79, 80, 81]. Here, we find that the optimal measurements can be achieved via anti-TAT-and-turn, regarded as the time-reversal process of TAT-and-turn. More importantly, the anti-TAT-and-turn can be realized by Floquet engineering without the need of flipping the sign of nonlinear interaction $\chi$.
With the generated non-Gaussian entangled state, one can encode the estimated phase along the best sensing direction by $\hat{R}_{\vec{n}_{\rm max}}(\phi)=e^{-i\phi\hat{J}_{\vec{n}_{\rm max}}}$. Based upon the Hamiltonian (9), adjusting the frequency of coupling field and the phase of Rabi frequency, $\delta\mapsto-\delta$ and $\alpha=0\mapsto\alpha=\pi/2$, one can obtain an effective anti-TAT-and-turn Hamiltonian $\hat{H}_{\rm{a-TATNT}}^{\alpha=0}=-\hat{H}_{\rm TATNT}^{\alpha=0}$. Thus the state before measurement can be written as
$\lvert\psi_{f}\rangle$ $=$ $\hat{U}_{2}\hat{R}_{\vec{n}}(\phi)\hat{U}_{1}\lvert\psi_{i}\rangle$ (7)
$=$ $e^{-i\int\hat{H}_{\rm FE}^{\alpha=\pi/2,-\delta}t}e^{-i\phi\hat{J}_{\vec{n}_{\rm max}}}e^{-i\int\hat{H}_{\rm FE}^{\alpha=0,\delta}t}\lvert\psi_{i}\rangle$
$\approx$ $e^{-i\hat{H}_{\rm{TATNT}}^{\alpha=\pi/2,-\delta}t}e^{-i\phi\hat{J}_{\vec{n}_{\rm max}}}e^{-i\hat{H}_{\rm TATNT}^{\alpha=0,\delta}t}\lvert\psi_{i}\rangle,$
where $\lvert\psi_{i}\rangle=\lvert\uparrow\rangle^{\otimes N}$, $\hat{H}_{\rm TATNT}^{\alpha=0,\delta}=\chi_{\rm eff}(\hat{J}_{y}^{2}-\hat{J}_{x}^{2})+\delta_{\rm eff}\hat{J}_{z}$ and $\hat{H}_{\rm{TATNT}}^{\alpha=\pi/2,-\delta}=-\hat{H}_{\rm{TATNT}}^{\alpha=0,\delta}=-\chi_{\rm eff}(\hat{J}_{y}^{2}-\hat{J}_{x}^{2})-\delta_{\rm eff}\hat{J}_{z}$. Therefore, to find out the optimal measurement, one can optimize the sensitivity over all possible directions through singular value decomposition
$\mathop{\max}_{\vec{m}}[\Delta\phi(\vec{n}_{\rm max},\vec{m})]^{-2}=\vec{n}_{\rm max}^{T}\mathcal{K}\vec{n}_{\rm max},$ (8)
where $\mathcal{K}=\mathcal{M}\mathcal{Q}^{-1}\mathcal{M}^{T}$ with two $3\times 3$ matrices $\mathcal{M}$ and $\mathcal{Q}$, whose elements are given as $\mathcal{M}_{\mu\nu}=i\langle[\hat{U}^{\dagger}_{1}\hat{J}_{\mu}\hat{U}_{1},\hat{\tilde{J}}_{\nu}]\rangle_{i}$ and $\mathcal{Q}_{\mu\nu}=\langle\{\hat{\tilde{J}}_{\mu},\hat{\tilde{J}}_{\nu}\}\rangle_{i}/2-\langle\hat{\tilde{J}}_{\mu}\rangle_{i}\langle\hat{\tilde{J}}_{\nu}\rangle_{i}$ for $\mu,\nu=x,y,z$ with $\hat{\tilde{J}}_{\mu(\nu)}=\hat{U}^{\dagger}_{1}\hat{R}_{\vec{n}_{\rm max}}^{\dagger}(\phi)\hat{U}^{\dagger}_{2}\hat{J}_{\mu(\nu)}\hat{U}_{2}\hat{R}_{\vec{n}_{\rm max}}(\phi)\hat{U}_{1}$ [94, 95]. The optimal measurement direction is $\vec{m}_{\rm max}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ with the normalization constant $\mathcal{N}_{c}$ determined by the Cauchy-Schwarz inequality [66]. Thus one can easily rotate the final state and measure the half-population difference $\hat{J}_{z}$ to infer $\phi$. Notably, the above calculations are valid for arbitrary phase $|\phi|\geq 0$, which generalize the result only for the zero phase $\phi=0$ [94, 95]. As shown in Fig. 3 (a), the nonlinear readout with Floquet-engineered TAT-and-turn achieves better measurement precision than those with TAT, OAT-and-turn and OAT. The measurement precision of our protocol can approach the corresponding QCRB, even for $t\lesssim t_{\rm opt}$.
Furthermore, an important capability of the Floquet-engineered anti-TAT-and-turn readout is its robustness against detection noise. In Fig. 3 (b), we show that our protocol is much more robust than using parity measurement. Here, we assume the detector possessing Gaussian noise and so that the measurement outcome is described by $\langle{\hat{J}}_{\vec{m}}\rangle_{\sigma}=\sum_{m=-N/2}^{N/2}P_{m}(\phi|\sigma)m$. Here $P_{m}(\phi|\sigma)=\sum_{n=-N/2}^{N/2}C_{n}e^{-(m-n)^{2}/2\sigma^{2}}P_{n}(\phi)$ represents the probability distribution under detection noise, $C_{n}=1/\sum_{m=-N/2}^{N/2}\mathrm{e}^{-(m-n)^{2}/2\sigma^{2}}$ is the normalization factor and $P_{n}(\phi)$ denotes the ideal probability distribution. Especially for $\sigma\geq 1$, the parity measurement becomes invalid and our protocol can still attain metrological gain beyond the SQL. Our Floquet-engineered anti-TAT-and-turn can be achieved by adjusting the periodic coupling without changing the interaction, which is feasible for practical entanglement-enhanced sensing.
Conclusions.— We use the continuous Floquet engineering with periodically-modulated Rabi frequency under appropriate detuning to achieve an effective TAT-and-turn dynamics. With semi-classical and quantum approaches, we analytically find the preparation timescale at large particle number is significantly shorter compared to the OAT, and even faster than using artificial three-body interactions [93]. Moreover, applying another Floquet-engineered TAT-and-turn dynamics, one can realize the time-reversal readout without flipping the sign of the nonlinear interaction. Our work provides an alternative way for fast generation and detection of large-particle-number non-Gaussian entangled states, which can be created in various systems [96, 97, 98, 99, 100, 101, 102, 103, 104] and has promising applications in quantum metrology and quantum information science.
Acknowledgements.
The authors would like to thank Dr. Wenjie Liu for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12025509) and the National Key Research and Development Program of China (Grant No. 2022YFA1404104).
Supplemental Material
I I. Derivation of Floquet-engineered TAT-and-turn Hamiltonian
By periodically modulating an external field coupled to an ensemble of identical Bose condensed two-level atoms, the Floquet Engineered (FE) Hamiltonian $\hat{H}_{\rm FE}$ is written as
$\hat{H}_{\rm FE}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega(t)\hat{J}_{\alpha},$ (9)
where $\Omega(t)=\Omega_{0}\cos(\omega t)$, and $\hat{J}_{\alpha}=\hat{J}_{x}\cos\alpha+\hat{J}_{y}\sin\alpha$. To perform the analytic calculation, we first transform from the Schrödinger picture to the interaction picture, rewriting $\hat{H}_{\rm FE}$ as a constant term, $\hat{H}_{0}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}$, plus a time-dependent term, $\hat{V}(t)=\Omega_{0}\cos(\omega t)\hat{J}_{\alpha}$. Applying the Baker-Campbell-Hausdorff (BCH) formula, we obtain $e^{i\gamma\hat{J}_{\alpha}}\hat{J}_{z}e^{-i\gamma\hat{J}_{\alpha}}=\cos\gamma\hat{J}_{z}+\sin\gamma\hat{J}_{\beta}$ with $\beta=\alpha+\pi/2$. In the interaction picture, we derive
$\hat{H}_{I}=$ $\hat{U}^{\dagger}\hat{H}_{\rm FE}\hat{U}-i\hat{U}^{\dagger}\partial_{t}\hat{U}$ (10)
$=$ $\hat{U}^{\dagger}\hat{H}_{0}\hat{U}$
[…]
$=$ $\chi(\hat{J}_{z}\cos\gamma+\hat{J}_{\beta}\sin\gamma)^{2}+\delta(\hat{J}_{z}\cos\gamma+\hat{J}_{\beta}\sin\gamma),$
[…]
According to the Euler formula $e^{i\gamma}=\cos{\gamma}+i\sin{\gamma}$ and the Jacobi-Anger expansion variant $e^{iz\sin\theta}=\mathcal{{J}}_{0}(|z|)+\sum_{n=1}^{\infty}\left[\mathcal{{J}}_{n}(z)e^{in\theta}+(-1)^{n}\mathcal{{J}}_{n}(z)e^{-in\theta}\right]$ , where $\mathcal{{J}}_{n}(z)$ is the $n$-th Bessel function of the first kind, we perform the Fourier decomposition and the Hamiltonian can be expressed as
$\hat{H}_{I}=\hat{H}_{0}^{I}+\sum_{n=1}^{\infty}(\hat{H}_{n}^{I}e^{in\omega t}+\hat{H}_{-n}^{I}e^{-in\omega t})$ (11)
where the $n$-order terms in the Hamiltonian are given by
$\hat{H}_{0}^{I}=$ $\frac{\chi}{2}[(1+L_{0})\hat{J}_{z}^{2}+(1-L_{0})\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z},$ (12)
$\hat{H}_{n}^{I}=$ $\frac{\chi}{4}L_{n}[(-1)^{n}\hat{J}_{1}^{2}+\hat{J}_{2}^{2}]+\frac{\delta}{2}K_{n}[(-1)^{n}\hat{J}_{1}+\hat{J}_{2}],$
$\hat{H}_{-n}^{I}=$ $\frac{\chi}{4}L_{n}[\hat{J}_{1}^{2}+(-1)^{n}\hat{J}_{2}^{2}]+\frac{\delta}{2}K_{n}[\hat{J}_{1}+(-1)^{n}\hat{J}_{2}],$
with $\hat{J}_{1,2}=\hat{J}_{z}\pm i\hat{J}_{\beta}$, $L_{n}=\mathcal{{J}}_{n}\left(2\Omega_{0}/\omega\right)$ and $K_{n}=\mathcal{{J}}_{n}\left(\Omega_{0}/\omega\right)$. When $n$ is odd (denoted by $p$), we have $H_{-p}^{I}=-H_{p}^{I}$, so that the commutation relations between Hamiltonians are
$[\hat{H}_{p}^{I},\hat{H}_{-p}^{I}]=0,$ (13)
$[\hat{H}_{p}^{I},\hat{H}_{0}^{I}]=a_{p}\chi^{2}\{\hat{J}_{z},\{\hat{J}_{z},\hat{J}_{\alpha}\}\}+b_{p}\chi^{2}\{\hat{J}_{\beta},\{\hat{J}_{\beta},\hat{J}_{\alpha}\}\}+c_{p}\chi\delta\{\hat{J}_{\alpha},\hat{J}_{z}\}+d_{p}\delta^{2}\hat{J}_{\alpha},$
$[\hat{H}_{-p}^{I},\hat{H}_{0}^{I}]=-[\hat{H}_{p}^{I,1},\hat{H}_{0}^{I}].$
with
$a_{p}=\frac{1}{4}L_{p}(L_{0}+1),~{}b_{p}=\frac{1}{4}L_{p}(L_{0}-1),~{}c_{p}=\frac{1}{2}(L_{p}K_{0}+L_{0}K_{p}+K_{p}),~{}d_{p}=K_{0}K_{p},$ (14)
where $[\hat{P},\hat{Q}]=\hat{P}\hat{Q}-\hat{Q}\hat{P}$ and $\{\hat{P},\hat{Q}\}=\hat{P}\hat{Q}+\hat{Q}\hat{P}$ are commutation and anti-commutation, respectively. In addition, when $n$ is even (denoted by $q$), we have $H_{-q}^{I}=H_{q}^{I}$, the commutation relations become
$=0,$ (15)
$[\hat{H}_{-q}^{I},\hat{H}_{0}^{I}]$ $=[\hat{H}_{q}^{I},\hat{H}_{0}^{I}].$
According to the Floquet-Magnus expansion [89, 90], the effective Floquet Hamiltonian becomes
$\hat{H}_{\mathrm{F}}^{\rm eff}$ $=\hat{H}_{0}^{I}+\sum_{n=1}^{\infty}(\frac{[\hat{H}_{n}^{I},\hat{H}_{-n}^{I}]}{n\omega}-\frac{[\hat{H}_{n}^{I},\hat{H}_{0}^{I}]}{n\omega}+\frac{[\hat{H}_{-n}^{I},\hat{H}_{0}^{I}]}{n\omega})+\mathcal{O}(\omega^{-1})$ (16)
$=\hat{H}_{0}^{I}-\sum_{p=1}^{\infty}\frac{2}{p\omega}[\hat{H}_{p}^{I},\hat{H}_{0}^{I}]+\mathcal{O}(\omega^{-1})$
$=\hat{H}_{0}^{I}-\frac{2\chi^{2}}{\omega}\{\hat{J}_{z},\{\hat{J}_{z},\hat{J}_{\alpha}\}\}\sum_{p=1}^{\infty}\frac{a_{p}}{p}-\frac{2\chi^{2}}{\omega}\{\hat{J}_{\beta},\{\hat{J}_{\beta},\hat{J}_{\alpha}\}\}\sum_{p=1}^{\infty}\frac{b_{p}}{p}-\frac{2\chi\delta}{\omega}\{\hat{J}_{\alpha},\hat{J}_{z}\}\sum_{p=1}^{\infty}\frac{c_{p}}{p}-\frac{2\delta^{2}}{\omega}\hat{J}_{\alpha}\sum_{p=1}^{\infty}\frac{d_{p}}{p}+\mathcal{O}(\omega^{-1}).$
It indicates that the $q$(even)-order terms make no contribution to the effective Floquet Hamiltonian. Further, in order to keep only $\hat{H}_{0}^{I}$ in the Hamiltonian (16), corresponding to drop the time-dependent terms in the Hamiltonian (11), we require $\omega$ to be sufficiently large $(\omega\gg\chi,\delta)$. Then the Hamiltonian becomes
$\hat{H}_{\rm F}^{\mathrm{eff}}\simeq\hat{H}_{0}^{I}=\frac{\chi}{2}[(1+L_{0})\hat{J}_{z}^{2}+(1-L_{0})\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z},$ (17)
which reveals that the external driving field leads to the twisting effect along both $z$ and $\beta$ directions. This can be interpreted intuitively as the rotation of spins perpendicular to the axis of OAT ($z$ axis) diverted from the twisting axis. Rewriting the Hamiltonian by adding a constant $-\frac{\chi}{2}(1+L_{0})\hat{J}^{2}$ ($\hat{J}^{2}=\hat{J}_{\alpha}^{2}+\hat{J}_{\beta}^{2}+\hat{J}_{z}^{2}$ is conserved during the dynamics), we obtain a mixture of an OAT Hamiltonian and a TAT Hamiltonian as
$\hat{H}_{\rm F}$ $=-\frac{\chi}{2}[(1+3L_{0})\hat{J}_{\alpha}^{2}+2L_{0}(\hat{J}_{\beta}^{2}-\hat{J}_{\alpha}^{2})]+K_{0}\delta\hat{J}_{z}$ (18)
$=-\frac{\chi}{2}[(1+L_{0})\hat{J}_{\alpha}^{2}+2L_{0}\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z}.$
When $L_{0}=-1/3$, corresponding to ${\Omega_{0}}/{\omega}\simeq 1.6262$, the Hamiltonian becomes effective TAT-and-turn Hamiltonian
$\hat{H}_{\rm TATNT}=\chi_{\rm eff}(\hat{J}_{\beta}^{2}-\hat{J}_{\alpha}^{2})+\delta_{\rm eff}\hat{J}_{z},$ (19)
where the effective nonlinear interaction strength $\chi_{\mathrm{eff}}=\chi/3$ and the effective detuning $\delta_{\mathrm{eff}}=\delta{K}_{0}\simeq 0.404\delta$. Further, by adjusting the frequency of Rabi frequency and the phase of external coupling field ($\alpha\mapsto\alpha+\pi/2$ and $\delta\mapsto-\delta$), one can obtain the anti-TAT-and-turn Hamiltonian
$\hat{H}_{\rm TATNT}^{\rm anti}=\chi_{\rm eff}(\hat{J}_{\alpha}^{2}-\hat{J}_{\beta}^{2})-\delta_{\rm eff}\hat{J}_{z},$ (20)
which can be applied for time-reversal interaction-based readout without flipping the nonlinear interaction.
II II. The maximum quantum Fisher information
In this section, we explain the method of diagonalizing the matrix to obtain the maximum QFI. In quantum metrology, the precision of estimating a parameter $\phi$ is bounded by the Quantum Cramér-Rao Bound (QCRB), given by
$\Delta\phi\geq\frac{1}{\sqrt{\nu F_{Q}}},$ (21)
where $\nu$ is the number of identical experiments, $F_{Q}$ is quantum Fisher information (QFI). For pure states, the QFI is written as
$F_{Q}=4(\Delta\hat{J}_{\vec{n}})^{2}=4(\langle\hat{J}^{2}_{\vec{n}}\rangle-\langle\hat{J}_{\vec{n}}\rangle^{2}),$ (22)
where $\hat{J}_{\vec{n}}$ is the generator for sensing with $e^{-i\hat{J}_{\vec{n}}\phi}$. For a given state, the optimal QFI $F_{Q}$ is determined by the optimal signal encoding direction. For obtaining the optimal signal encoding direction, one can decompose the $\hat{J}_{\vec{n}}=n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z}$ with $n_{x}^{2}+n_{y}^{2}+n_{z}^{2}=1$. Therefore the QFI is written as
$F_{Q}=$ $4\langle(n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z})^{2}\rangle-4\langle n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z}\rangle^{2}$ (23)
$=$ $n_{x}n_{x}(\langle 2\{\hat{J}_{x},\hat{J}_{x}\}\rangle-4\langle\hat{J}_{x}\rangle\langle\hat{J}_{x}\rangle)+n_{x}n_{y}(2\langle\{\hat{J}_{x},\hat{J}_{y}\}\rangle-4\langle\hat{J}_{x}\rangle\langle\hat{J}_{y}\rangle)+n_{x}n_{z}(2\langle\{\hat{J}_{x},\hat{J}_{z}\}\rangle-4\langle\hat{J}_{x}\rangle\langle\hat{J}_{z}\rangle)$
$+n_{y}n_{x}(2\langle\{\hat{J}_{y},\hat{J}_{x}\}\rangle-4\langle\hat{J}_{y}\rangle\langle\hat{J}_{x}\rangle)+n_{y}n_{y}(2\langle\{\hat{J}_{y},\hat{J}_{y}\}\rangle-4\langle\hat{J}_{y}\rangle\langle\hat{J}_{y}\rangle)+n_{y}n_{z}(2\langle\{\hat{J}_{y},\hat{J}_{z}\}\rangle-4\langle\hat{J}_{y}\rangle\langle\hat{J}_{z}\rangle)$
$+n_{z}n_{x}(2\langle\{\hat{J}_{z},\hat{J}_{x}\}\rangle-4\langle\hat{J}_{z}\rangle\langle\hat{J}_{x}\rangle)+n_{z}n_{y}(2\langle\{\hat{J}_{z},\hat{J}_{y}\}\rangle-4\langle\hat{J}_{z}\rangle\langle\hat{J}_{y}\rangle)+n_{z}n_{z}(2\langle\{\hat{J}_{z},\hat{J}_{z}\}\rangle-4\langle\hat{J}_{z}\rangle\langle\hat{J}_{z}\rangle).$
To find the maximum QFI, one needs to optimize over all possible directions $\vec{n}$. This optimization problem can be recast in terms of a matrix diagonalization problem, where the QFI can be represented in matrix form:
$F_{Q}=\begin{pmatrix}n_{x}&n_{y}&n_{z}\\ \end{pmatrix}\begin{pmatrix}\mathcal{F}_{xx}&\mathcal{F}_{xy}&\mathcal{F}_{xz}\\ \mathcal{F}_{yx}&\mathcal{F}_{yy}&\mathcal{F}_{yz}\\ \mathcal{F}_{zx}&\mathcal{F}_{zy}&\mathcal{F}_{zz}\\ \end{pmatrix}\begin{pmatrix}n_{x}\\ n_{y}\\ n_{z}\end{pmatrix},$ (24)
with the Quantum Fisher Information Matrix (QFIM) $\mathcal{F}$ defined as [91]:
$\mathcal{F}=\begin{pmatrix}\mathcal{F}_{xx}&\mathcal{F}_{xy}&\mathcal{F}_{xz}\\ \mathcal{F}_{yx}&\mathcal{F}_{yy}&\mathcal{F}_{yz}\\ \mathcal{F}_{zx}&\mathcal{F}_{zy}&\mathcal{F}_{zz}\\ \end{pmatrix},$ (25)
where the elements are given by
$\mathcal{F}_{\mu\nu}=2\langle\{\hat{J}_{\mu},\hat{J}_{\nu}\}\rangle-4\langle\hat{J}_{\mu}\rangle\langle\hat{J}_{\nu}\rangle,$ (26)
with $\mu,\nu=x,y,z$. Therefore, the maximum QFI $F_{Q}^{\mathrm{max}}$ is the largest eigenvalue of QFIM,
$\mathcal{F}\vec{n}_{\mathrm{max}}=\lambda_{\mathrm{max}}\vec{n}_{\mathrm{max}}=F_{Q}^{\mathrm{max}}\vec{n}_{\mathrm{max}},$ (27)
where the optimal signal encoding direction vector $\vec{n}_{\mathrm{max}}=(n_{x}^{\mathrm{max}},n_{y}^{\mathrm{max}},n_{z}^{\mathrm{max}})^{\mathrm{T}}$, and the responding optimal generator is $\hat{J}_{\vec{n}_{\mathrm{max}}}=n_{x}^{\mathrm{max}}\hat{J}_{x}+n_{y}^{\mathrm{max}}\hat{J}_{y}+n_{z}^{\mathrm{max}}\hat{J}_{z}$.
III III. Semi-classical treatment of TAT-and-turn dynamics
We use the semi-classical treatment, introduced in Ref. [92], to analyze the metrological properties of TAT-and-turn dynamics. Utilizing the commutations $[\hat{J}_{\alpha},\hat{J}_{\beta}]=i\epsilon_{\alpha\beta z}\hat{J}_{z}$ or its variants, we can describe the time-evolution of Hamiltonian (19) by the Heisenberg equations of motion, $d\hat{J}_{\kappa}/dt=i[\hat{H},\hat{J}_{\kappa}]$ where $\kappa=\alpha,\beta,z$ are the components of the collective spin. They are
$\frac{d\hat{J}_{\alpha}}{dt}$ $=i[\hat{H}_{\mathrm{TATNT}},\hat{J}_{\alpha}]=\chi_{\mathrm{eff}}(\hat{J}_{\beta}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{\beta})-\delta_{\mathrm{eff}}\hat{J}_{\beta},$ (28)
$\frac{d\hat{J}_{\beta}}{dt}$ $=i[\hat{H}_{\mathrm{TATNT}},\hat{J}_{\beta}]=\chi_{\mathrm{eff}}(\hat{J}_{\alpha}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{\alpha})+\delta_{\mathrm{eff}}\hat{J}_{\alpha},$
$\frac{d\hat{J}_{z}}{dt}$ $=i[\hat{H}_{\mathrm{TATNT}},\hat{J}_{z}]=-2\chi_{\mathrm{eff}}(\hat{J}_{\alpha}\hat{J}_{\beta}+\hat{J}_{\beta}\hat{J}_{\alpha}).$
In the thermodynamic limit, $J=N/2\to\infty$, above Eqs.(28) lead to the phase-space flow of the classical variables $\mathbf{R}=(A,B,Z)=$ $(\langle\hat{{J}}_{\alpha}\rangle,\langle\hat{{J}}_{\beta}\rangle,\langle\hat{{J}}_{z}\rangle)/J$ with $A^{2}+B^{2}+Z^{2}=1$, which after neglecting correlations $\langle\hat{P}\hat{Q}\rangle=\langle\hat{P}\rangle\langle\hat{Q}\rangle$, is given by
$\frac{dA}{dt}$ $=N\chi_{\mathrm{eff}}BZ-\delta_{\mathrm{eff}}B,$ (29)
$\frac{dB}{dt}$ $=N\chi_{\mathrm{eff}}AZ+\delta_{\mathrm{eff}}A,$
$\frac{dZ}{dt}$ $=-2N\chi_{\mathrm{eff}}AB.$
Likewise, the classical phase-space trajectories of nonlinear or linear modulation are plotted in Fig.1 (c) of the main text.
The direction of the initial spin coherent state of the non-adiabatic evolution of the Hamiltonian (19) is determined by fixed points of a phase-space flow with a trivial evolution. By solving $d\mathbf{R}/dt=0$, the phase-space flow in Eqs. (29) has six different fixed points depending on the system parameters (we consider the conditions about $\chi\neq 0$, and $|\delta_{\mathrm{eff}}/(N\chi_{\mathrm{eff}})|<1$) are given by
$(A,B,Z)$ $=(0,0,\pm 1),$ (30)
$(A,B,Z)$ $=\left(0,\pm\sqrt{1-\left(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}\right)^{2}},\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}\right),$
$(A,B,Z)$ =(±$=\left(\pm\sqrt{1-\left(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}\right)^{2}},0,-\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}\right).$
For the fixed points, the stability analysis comes from the diagonalization of the Jacobi matrix, ${M}[\mathbf{R}]=(\partial/\partial\mathbf{R})(d\mathbf{R}/dt)$. In Eqs. (29), this matrix of the phase-space flow is given by
${M}[\mathbf{R}]=\begin{pmatrix}0&&N\chi_{\mathrm{eff}}Z-\delta_{\mathrm{eff}}&&N\chi_{\mathrm{eff}}B\\ N\chi_{\mathrm{eff}}Z+\delta_{\mathrm{eff}}&&0&&N\chi_{\mathrm{eff}}A\\ -2N\chi_{\mathrm{eff}}B&&-2N\chi_{\mathrm{eff}}A&&0\end{pmatrix}.$ (31)
The analysis shows that two fixed points $(A,B,Z)=(0,0,\pm 1)$ that always pointing toward the ±$\pm z$-axis are saddle points and the other four points are stable points. The Jacobi matrix of the saddle points is
${M}[\mathbf{R}]=\begin{pmatrix}0&&\pm N\chi_{\mathrm{eff}}-\delta_{\mathrm{eff}}&&0\\ \pm N\chi_{\mathrm{eff}}+\delta_{\mathrm{eff}}&&0&&0\\ 0&&0&&0\end{pmatrix}.$ (32)
Figure 4: The maximum QFI $F_{Q}^{\mathrm{max}}$ obtained via TAT-and-turn dynamics varying in short evolution time with different $\delta$ for particle number $N=100$. The white dashed line represents the maximum varying in time, and indicates that the QFI has the fastest growth rate within short evolution time when $\delta_{\mathrm{eff}}=0$.
In general, the separatrix branches emerge from this saddle point, which yield the classical evolution trajectory of the Hamiltonian. By evaluating the eigenvalues of the Jacobi matrix in Eq. (32) at this saddle, one obtains its local Lyapunov exponent
$\Lambda_{\mathrm{sd}}^{\mathrm{TATNT}}=N\chi_{\mathrm{eff}}\sqrt{1-\biggl{(}\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}\biggr{)}^{2}}.$ (33)
It dictates the exponential rate at which points move away from the saddle point, i.e., the optimal entanglement generation parameters at the initial moment. This rate reaches maximum when $\delta_{\mathrm{eff}}=0$, which is equivalent to the TAT Hamiltonian and verified to generate metrologically useful states within a short time in Fig. 4, and the corresponding Lyapunov exponent is given by $\Lambda_{\mathrm{sd}}^{\mathrm{TAT}}=N\chi_{\mathrm{eff}}$.
Besides, we can obtain the energy density of the TAT-and-turn Hamiltonian by computing $\langle\hat{H}_{\mathrm{TATNT}}\rangle/J$ in the thermodynamic limit. After neglecting correlations as before and in terms of the classical variables $\mathbf{R}=(A,B,Z)$, this energy density in classical phase space becomes
$E(A,B,Z;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=\frac{\langle\hat{H}_{\mathrm{TATNT}}\rangle}{J}=\delta_{\mathrm{eff}}Z-\frac{N\chi_{\mathrm{eff}}}{2}(A^{2}-B^{2}).$ (34)
Since the conservation of energy guarantees that points on the separatrix have the same energy as the one of the saddle, the separatrix equation of two saddles can be constructed by $E(A,B,Z;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=E(0,0,\pm 1;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=\pm\delta_{\mathrm{eff}}$. Due to $A^{2}+B^{2}+Z^{2}=1$, the above formula can be rewritten as
$E(A,B,Z;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=\delta_{\mathrm{eff}}Z-\frac{N\chi_{\mathrm{eff}}}{2}(2A^{2}+Z^{2}-1)=\pm\delta_{\mathrm{eff}},$ (35)
or
$E(A,B,Z;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=\delta_{\mathrm{eff}}Z-\frac{N\chi_{\mathrm{eff}}}{2}(1-Z^{2}-2B^{2})=\pm\delta_{\mathrm{eff}}.$ (36)
The relationship between classical space components is obtained as
$A^{2}=\frac{1}{2}(1\mp Z)(1\pm Z\mp\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}),$ (37)
and
$B^{2}=\frac{1}{2}(1\mp Z)(1\pm Z\pm\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}).$ (38)
We can further estimate the global optimal parameter $\delta_{\rm opt}$ and the explicit expression for the time evolution to reach maximum QFI using the semi-classical description. We find that the TAT-and-turn Hamiltonian can generate GHZ-like states, whose variance is stretched along $\alpha$-axis or $\beta$-axis on the $x-y$ plane. For calculation clarity, we start from the initial position $(0,0,1)$, when $\delta_{\mathrm{eff}}>0$ (default $\chi>0$), we take $\hat{J}_{\beta}$ as the generator and obtain
$F_{Q}=4\left(\Delta\hat{J}_{\beta}\right)^{2}=N^{2}B^{2}=\frac{N^{2}}{2}(1-Z)(1+Z+\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}).$ (39)
When $Z=\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}$, the above Eq. (39) reaches its maximum about variable $Z$, and is converted into
$F_{Q}=\frac{N^{2}}{2}(1-\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}\leq(F_{Q})_{\mathrm{max}}\leq N^{2}.$ (40)
One can obtain the maximum QFI $(F_{Q})_{\mathrm{max}}$ when
${\delta_{\rm opt}}/({N\chi_{\mathrm{eff}}})=\sqrt{2}-1\approx 0.414.$ (41)
And for $\delta_{\mathrm{eff}}<0$, the expression of QFI is $F_{Q}=4\left(\Delta\hat{J}_{\alpha}\right)^{2}=N^{2}A^{2}\leq N^{2}$ with the generator being $\hat{J}_{\alpha}$, the corresponding optimal detuning is ${\delta_{\rm opt}}/({N\chi_{\mathrm{eff}}})=1-\sqrt{2}\approx-0.414$.
We consider $\delta_{\mathrm{eff}}>0$ (the calculation process for $\delta_{\mathrm{eff}}<0$ is the same as below), within the time range $0\leq t\leq t_{\rm opt}$ of reaching the maximum QFI. We find that $AB=\frac{1}{2}(1-Z)\sqrt{(1+Z)^{2}-(\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}>0$, and obtain a differential equation that depends only on variable $Z$, which is given by
$\frac{dZ}{dt}=-2N\chi_{\mathrm{eff}}AB=-N\chi_{\mathrm{eff}}(1-Z)\sqrt{(1+Z)^{2}-(\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}},$ (42)
and the desired timescale is given by the integration
$N\chi_{\mathrm{eff}}t_{\rm opt}=-\int_{Z(0)}^{Z(t_{\rm opt})}\frac{dZ}{(1-Z)\sqrt{(1+Z)^{2}-(\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}$ $=-\frac{f(Z)}{\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\bigg{|}_{Z(0)}^{Z(t_{\rm opt})}$ (43)
$\approx-\frac{f(Z)}{\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\bigg{|}_{\sqrt{1-\frac{1}{2N}}}^{\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1}$
$\approx\frac{\ln\left\{\frac{16\left[1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}\right]}{1-\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\right\}-2\operatorname{arctanh}\left[\frac{1-\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}}{\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\right]+\ln{N}}{2\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}},$
with
$f(Z)=\operatorname{arctanh}\left[\frac{1-Z+\sqrt{\left(1+Z\right)^{2}-(\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}{2\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\right].$ (44)
This result is obtained via the property of uncertainty patch. The lower bound $\sqrt{1-\frac{1}{2N}}$ of the integral comes from the separatrix of the uncertainty patch, which determines the initial stretching position, and the integral upper bound $\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1$ is the critical point of variance change on the separatrix. The Taylor expansion used above is
$\operatorname{arctanh}\left[\frac{1+\sqrt{\left(1+\sqrt{1-\frac{1}{2N}}\right)^{2}-4(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}-\sqrt{1-\frac{1}{2N}}}{2\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\right]=\frac{1}{2}\ln\left\{\frac{16\left[1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}\right]}{1-\sqrt{1-(\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}}\right\}+\frac{1}{2}\ln{N}+\mathcal{O}\left(\frac{1}{N}\right).$ (45)
When $\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}=\sqrt{2}-1$, $N\chi_{\mathrm{eff}}t_{\rm opt}\simeq{1.9+0.55\ln{N}}$, therefore
${\chi_{\rm eff}}t_{\rm opt}\simeq\frac{1.9+0.55\ln{N}}{N}.$ (46)
When $0\leq t\leq t_{\rm opt}$, we can determine $Z(t)$ by solving the unique root of the following equation:
$f[Z(t)]-f[Z(0)]+\sqrt{{(N\chi_{\mathrm{eff}}})^{2}-\delta_{\mathrm{eff}}^{2}}t=0,$ (47)
Due to the special symmetry of the phase space of the TAT-and-turn Hamiltonian, the same result can be obtained for the another saddle point $(0,0,-1)$, the calculation process is similar to the above.
IV IV. Heisenberg scaling and comparison with spin cat states
Figure 5: (a) The probability distributions of non-Gaussian state generated by Floquet-engineered TAT-and-turn dynamics and its corresponding spin cat state with $\theta=\arccos{\overline{Z}}$ for particle number $N=100$. (b) The QFI of non-Gaussian states generated by Floquet-engineered TAT-and-turn dynamics and their corresponding spin cat states with $\theta=\arccos{\overline{Z}}$ versus different particle number $N$.
In the main text, we have shown that the non-Gaussian states generated by Floquet-engineered TAT-and-turn dynamics have metrological properties of Heisenberg scaling, which are similar to the spin cat states. In this section, we further analyze the QFI scaling of the non-Gaussian states with optimal QFI generated using Floquet-engineered TAT-and-turn dynamic by approximating the generated non-Gaussian entangled states as spin cat states. Compared with using the semi-classical method of taking the critical point $Z(t_{\rm opt})\approx\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1$ through Eq. (39) in the previous section, the obtained results are closer to the numerical results using full quantum treatment.
We consider the non-Gaussian state $\lvert\psi\rangle_{\rm NGS}$ generated by Floquet-engineered TAT-and-turn dynamic starting from the initial state $\lvert\uparrow\rangle^{\otimes N}$ with $\alpha=0$, the time in Eq. (46) and the critical detuning in Eq. (41), whose two main parts are symmetrically distribution near the pole of $y$-axis. To unify the orientation with a spin-cat state consisting of two SCSs with the same azimuthal angle $\varphi$ and polar angle symmetric about $\vartheta=\pi/2$, the non-Gaussian state $\lvert\psi\rangle_{\rm NGS}$ need to be rotated to the $z$-axis via $e^{-i\hat{J}_{x}\pi/2}$. The expectation of $\hat{J}_{z}$ for the non-Gaussian state $\langle\hat{J}_{z}\rangle_{\rm NGS}^{{}^{\prime}}=0$ with $\lvert\psi\rangle_{\rm NGS}^{{}^{\prime}}=e^{-i\hat{J}_{x}\pi/2}\lvert\psi\rangle_{\rm NGS}$, and the probability distribution is mainly concentrated around the two ends of the $z$-axis. Thus the symmetrical non-Gaussian entangle state can be approximately regarded as a macroscopic superposition spin coherent state (MSSCS) like a spin cat state [48], which is given by
$\lvert\psi\rangle_{\rm NGS}^{{}^{\prime}}\sim|\Psi(\vartheta)\rangle_{\mathrm{CAT}}\approx\frac{1}{\sqrt{2}}(|\vartheta,\varphi\rangle_{\rm SCS}+|\pi-\vartheta,\varphi\rangle_{\rm SCS}),$ (48)
where $\vartheta=\arccos\overline{Z}$ and the normalized average spin length $\overline{Z}=\frac{2}{N}\sum_{m=-N/2}^{N/2}|m|P_{m}=\frac{2}{N}\sum_{m=-N/2}^{N/2}|m||\langle m|\psi\rangle^{{}^{\prime}}_{\rm{NGS}}|^{2}$ with the eigenstate $|m\rangle$ of $\hat{J}_{z}$. It approximates the non-Gaussian state as a spin cat state with the same normalized average spin length $\overline{Z}$. The probability distribution of this non-Gaussian state and its corresponding spin cat state are shown in Fig. 5 (a). In this case, the QFI of this non-Gaussian state is approximately written as
$F_{Q}^{\rm NGS}\sim F_{Q}^{\rm CAT}\approx N^{2}\cos^{2}\vartheta=N^{2}{\overline{Z}}^{2},$ (49)
Further, we use Eq. (49) to calculate the QFI of non-Gaussian states generated by TAT-and-turn dynamics for different particle number $N$ and find that the results are in agreement with the quantum results, as shown in Fig. 5 (b), illustrating that the best achievable QFI of Floquet-engineered TAT-and-turn exhibits the Heisenberg-limited scaling $F_{Q}^{\rm max}\propto N^{2}$, which is consistent with the QCRB of a spin cat state.
V V. Optimal first-order measurement
For a final state after given dynamics with interaction-based readout or half-population difference readout, the achievable measurement precision depends both on the directions to encode signal into probe state and to detect the final state. The signal encoding operator $\hat{J}_{\vec{n}}$ determines the QFI of a given probe state, thus setting a bound (i.e., QCRB) on the ultimate measurement precision. The detection using observable operator $\hat{J}_{\vec{m}}$ determines how closely we can approach the QCRB from a given measurement direction. Optimization in both directions $\vec{n}$ and $\vec{m}$ is of great significance to improving measurement precision. The method to optimize the measurement direction with a given sensing direction is shown as follows.
Figure 6: The interaction-based readout (IBR) at $\phi=1/1000$ for (a) OAT, (b) OAT-and-turn, (c) TAT, and (d) TAT-and-turn with particles number $N=100$ at evolution time $t=t_{1}=t_{2}$. (e) The attainable measurement precision of TAT-and-turn at $t=0.12$ via IBR for $\phi$ approaching 0.
In the general process of interaction-based readout, for an arbitrary observable operator $\hat{J}_{\vec{m}}$, we transform the Schrödinger picture into the Heisenberg picture, i.e.
$\hat{\tilde{J}}_{\vec{m}}$ $=\hat{U}^{\dagger}_{1}\hat{R}_{\vec{n}}^{\dagger}(\phi)\hat{U}^{\dagger}_{2}\hat{J}_{\vec{m}}\hat{U}_{2}\hat{R}_{\vec{n}}(\phi)\hat{U}_{1}$ (50)
$=\hat{U}^{\dagger}_{1}e^{i\phi\hat{J}_{\vec{n}}}\hat{U}_{1}\hat{U}^{\dagger}_{1}\hat{U}^{\dagger}_{2}\hat{J}_{\vec{m}}\hat{U}_{2}\hat{U}_{1}\hat{U}^{\dagger}_{1}e^{-i\phi\hat{J}_{\vec{n}}}\hat{U}_{1}$
$=e^{i\phi\hat{U}^{\dagger}_{1}\hat{J}_{\vec{n}}\hat{U}_{1}}\hat{U}^{\dagger}_{1}\hat{U}^{\dagger}_{2}\hat{J}_{\vec{m}}\hat{U}_{2}\hat{U}_{1}e^{-i\phi\hat{U}^{\dagger}_{1}\hat{J}_{\vec{n}}\hat{U}_{1}},$
where $\hat{R}_{\vec{n}}(\phi)=e^{-i\hat{J}_{\vec{n}}\phi}$ for encoding signal, $\hat{U}_{1}=e^{-i\hat{H}_{1}t_{1}}$ for generating entangled states, and $\hat{U}_{2}=e^{-i\hat{H}_{2}t_{2}}$ for signal amplification through disentanglement, typically time-reversal IBR with $\hat{H}_{2}=-\hat{H}_{1}$ and $t_{2}=t_{1}$. When half-population difference readout is performed directly instead of interaction-based readout, the Hamiltonian $\hat{H}_{2}=0$. Therefore we can calculate measurement response and express it in matrix form similar to Eq. (24) or Ref. [95], which is given by
$\frac{\partial\langle\hat{\tilde{J}}_{\vec{m}}\rangle}{\partial\phi}=i\langle[\hat{U}^{\dagger}_{1}\hat{J}_{\vec{n}}\hat{U}_{1},\hat{\tilde{J}}_{\vec{m}}]\rangle=\vec{n}^{T}\mathcal{M}\vec{m},$ (51)
where $\mathcal{M}$ is a $3\times 3$ matrix with elements $\mathcal{M}_{\mu\nu}=i\langle[\hat{U}^{\dagger}_{1}\hat{J}_{\mu}\hat{U}_{1},\hat{\tilde{J}}_{\nu}]\rangle(\mu,\nu=x,y,z)$. Similarly, the measurement uncertainty can be expressed as
$(\Delta\hat{\tilde{J}}_{\vec{m}})^{2}=\langle\hat{\tilde{J}}_{\vec{m}}^{2}\rangle-\langle\hat{\tilde{J}}_{\vec{m}}\rangle^{2}={\vec{m}}^{T}\mathcal{Q}{\vec{m}},$ (52)
where $\mathcal{Q}$ represents the $3\times 3$ covariance matrix with elements $\mathcal{Q}_{\mu\nu}=\langle\{\hat{\tilde{J}}_{\mu},\hat{\tilde{J}}_{\nu}\}\rangle/2-\langle\hat{\tilde{J}}_{\mu}\rangle\langle\hat{\tilde{J}}_{\nu}\rangle$. Based on the error propagation formula, the achievable measurement sensitivity becomes
$[\Delta\phi(\vec{n},\vec{m})]^{-2}=\frac{|\partial\langle\hat{\tilde{J}}_{\vec{m}}\rangle/\partial\phi|^{2}}{(\Delta\hat{\tilde{J}}_{\vec{m}})^{2}}=\frac{(\vec{n}^{T}\mathcal{M}\vec{m})^{2}}{\vec{m}^{T}\mathcal{Q}\vec{m}}.$ (53)
To find the optimal measurement direction, one can using the Cauchy-Schwartz inequality $(\vec{u}^{T}\vec{v})^{2}\leq\vec{u}^{T}\vec{u}\vec{v}^{T}\vec{v}$, the measurement sensitivity satisfies
$[\Delta\phi(\vec{n},\vec{m})]^{-2}\leq\mathop{\max}_{\vec{m}}[\Delta\phi(\vec{n},\vec{m})]^{-2}=\vec{n}^{T}\mathcal{K}\vec{n},$ (54)
in which $\mathcal{K}=\mathcal{M}\mathcal{Q}^{-1}\mathcal{M}^{T}$ depends on the selection of $\vec{n}$, and the equality is achieved at $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ with $\mathcal{N}_{c}$ the normalization constant, a given sensing operator $\hat{J}_{\vec{n}}$ for the general case including $\phi=0$ and $\phi\neq 0$.
Thus, after applying the optimal detection direction, i.e. $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$, the measurement precision only determined by the encoding signal direction $\vec{n}$. In this work, we choose the sensing direction $\vec{n}_{\rm max}$ corresponding to the maximum QFI value of the entangled state to encode the signal. Thus combining Eq. (27), Eq. (53) and Eq. (54), we arrive
$\mathop{\max}_{\vec{m}}[\Delta\phi(\vec{n}_{\rm max},\vec{m})]^{-2}=\vec{n}_{\rm max}^{T}\mathcal{K}\vec{n}_{\rm max}.$ (55)
Note that the phase $\phi$ cannot be equal to 00, otherwise the matrix $\mathcal{K}$ will lose the direction $\vec{n}_{\rm max}$ and be restricted to the plane perpendicular to the polarization direction.
As shown in Fig. 6 (a-d), we analyze the measurement precision of several common nonlinear Hamiltonians such as OAT, OAT-and-turn, TAT, and our TAT-and-turn under time-reversal IBR using above method. The results show that the method we proposed is close to the QCRB with excellent performance of measurement precision. The entangled state generated in a short time is a spin squeezed state, and after a longer time, the generated state becomes a non-Gaussian state. During the process, the measurement precision can approach the QCRB. Especially for OAT, this method can saturate QCRB in most of the time range, which achieves ideal measurement precision. From the Fig. 6 (e), we find that the measurement precision gradually converges as $\phi$ approaches 00, which is conducive to high-precision measurement of tiny signals.
VI VI. Experimental feasibility of Floquet-engineered TAT-and-turn dynamics
In an interacting atomic ensemble driven by an external coupling field, the system obey the Hamiltonian $\hat{H}=\chi\hat{J}_{z}^{2}+\omega_{0}\hat{J}_{z}+\Omega(t)[e^{i(\omega_{c}t+\alpha)}+e^{-i(\omega_{c}t+\alpha)}]\hat{J}_{x}$, where $\omega_{0}$ is the atomic transition frequency, $\omega_{c}$ is the frequency of the coupling field, $\alpha$ is the phase of coupling field, and $\Omega(t)$ is the Rabi frequency. To realize the Hamiltonian (9), we choose to modulate the Rabi frequency $\Omega(t)=2\Omega_{0}\cos(\omega t)$, i.e.
$\hat{H}=\chi\hat{J}_{z}^{2}+\omega_{0}\hat{J}_{z}+2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)\hat{J}_{x}$ (56)
In the rotating-frame with rotating-wave approximation (RWA), let $\hat{U}=e^{-i\omega_{c}t\hat{J}_{z}}$, the Hamiltonian becomes
$\hat{H}_{R}$ $=\hat{U}^{\dagger}\hat{H}\hat{U}-i\hat{U}^{\dagger}\frac{\partial\hat{U}}{\partial t}$ (57)
$=\chi\hat{J}_{z}^{2}+\omega_{0}\hat{J}_{z}+e^{i\omega_{c}t\hat{J}_{z}}\left[2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)\hat{J}_{x}\right]e^{-i\omega_{c}t\hat{J}_{z}}-\omega_{c}\hat{J}_{z}$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)e^{i\omega_{c}t\hat{J}_{z}}\hat{J}_{x}e^{-i\omega_{c}t\hat{J}_{z}}$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)\left[\cos(\omega_{c}t)\hat{J}_{x}-\sin(\omega_{c}t)\hat{J}_{y}\right]$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos\left(\omega t\right)\left\{\left[\cos(2\omega_{c}t+\alpha)+\cos\alpha\right]\hat{J}_{x}-\left[\sin(2\omega_{c}t+\alpha)-\sin\alpha\right]\hat{J}_{y}\right\}$
$\approx\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos\left(\omega t\right)\left(\hat{J}_{x}\cos\alpha+\hat{J}_{y}\sin\alpha\right)$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos\left(\omega t\right)\hat{J}_{\alpha},$
where the detuning $\delta=\omega_{0}-\omega_{c}$, and it shows that the Hamiltonian (9) we designed is feasible under near-resonance conditions.
Figure 7: Effects of driving frequency on Floquet-engineered TAT-and-turn for (a) the maximum QFI and (b) the attainable measurement precision via interaction-based readout. Here, $N=100$ and $\phi=1/1000$ are chosen for simulation.
In addition, we also verified that this Floquet-engineered Hamiltonian is equivalent to the TAT-and-turn Hamiltonian under certain conditions. From Fig. 7 we find that when the frequency is large enough, the Floquet-engineered Hamiltonian has the same metrological potential (e.g., QFI) and measurement precision as the ideal TAT-and-turn Hamiltonian. Moreover, under not so high-frequency driving, the measurement precision of interaction-based readout is close to the prediction, which is conducive to experimental implementation.
At last, in real experiments or application scenarios, both the sensing and measurement operators will select $\hat{J}_{z}$, the former for free evolution (phase accumulation) and the latter for half-population readout. We can equivalently perform the desired $\hat{J}_{\vec{n}}$ and $\hat{J}_{\vec{m}}$ with the pulses in the $x-y$ plane. To achieve an equivalent generator $\hat{J}_{\vec{n}}$, assuming the pulses are along $x$-axis and $y$-axis respectively, four pulses are applied before and after the free evolution. Thus we have
$e^{i\vartheta_{n}\hat{J}_{x}}e^{i\varphi_{n}\hat{J}_{y}}e^{-i\phi\hat{J}_{z}}e^{-i\varphi_{n}\hat{J}_{y}}e^{-i\vartheta_{n}\hat{J}_{x}}$ $=e^{i\vartheta_{n}\hat{J}_{x}}e^{-i\phi(\hat{J}_{z}\cos\varphi_{n}-\hat{J}_{x}\sin\varphi_{n})}e^{-i\vartheta_{n}\hat{J}_{x}}$ (58)
$=e^{-i\phi[(\hat{J}_{z}\cos\vartheta_{n}+\hat{J}_{y}\sin\vartheta_{n})\cos\varphi_{n}-\hat{J}_{x}\sin\varphi_{n}]}$
$=e^{-i\phi(-\hat{J}_{x}\sin\varphi_{n}+\hat{J}_{y}\cos\vartheta_{n}\sin\vartheta_{n}+\hat{J}_{z}\cos\varphi_{n}\cos\vartheta_{n})}.$
When setting $\vartheta_{n}=\arctan{(n_{y}/n_{z})}$, $\varphi_{n}=-\arcsin{n_{x}}$, the above formula is equivalent to $e^{-i\phi\hat{J}_{\vec{n}}}=e^{-i\phi(n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z})}$. Similarly, for half-population readout, one requires two pulses applied to the final state, and we can obtain the following equation
$\langle\hat{J}_{\vec{m}}\rangle_{f}=\langle e^{i\vartheta_{m}\hat{J}_{x}}e^{i\varphi_{m}\hat{J}_{y}}\hat{J}_{z}e^{-i\varphi_{m}\hat{J}_{y}}e^{-i\vartheta_{m}\hat{J}_{x}}\rangle_{f},$ (59)
where $\vartheta_{m}=\arctan{(m_{y}/m_{z})}$, $\varphi_{m}=-\arcsin{m_{x}}$. And the required sensing direction $\vec{n}$ can be obtained by constructing QFIM (25), the corresponding optimal measurement direction $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ in time-reversal IBR.